2 cyclists leave from home and travel in straight lines that form an angle of 27 degrees and 11 minutes. The 1st travels 45.3 miles and the 2nd travels 53.7 miles. When they stop how far apart are they?
**more info if needed
of triangle ABC, C is the hypotenuse. ** not needed.

Answer by stanbon(48558)

2 cyclists leave from home and travel in straight lines that form an angle of 27 degrees and 11 minutes. The 1st travels 45.3 miles and the 2nd travels 53.7 miles. When they stop how far apart are they?
------------------------
Draw the picture including a line segment joining the two cyclists; call that
segment "a".
------------------
Using the Law of Cosines:
a^2 = 45.3^2 + 53.7^2-2*45.3*53.7Cos(27 11')
a^2 = 607.9369... miles
a = 24.66 miles
====================
Cheers,
Stan H.